1. Write an equation of the line that passes through (10,-4) and is perpendicular to the line whose equation is y = 2/5x - 8.



A.
y = 2/5x - 8



B.
y = 2/5x + 58/5



C.
y = -5/2x + 21



D.
y = -5/2x - 29



Answer by Alan3354(69443)

Write an equation of the line that passes through (10,-4) and is perpendicular to the line whose equation is y = 2/5x - 8.
------------------
The slope, m, of the given eqn is 2/5.
The slope of lines perpendicular is the negative inverse, = -5/2.
Use y = mx+b and the point to find b.
-4 = (-5/2)*10 + b
b = 21
------
y = (-5/2)x + 21
----------------
All that space and the choices are not needed.